# Characterizing pedestrian contact interaction trajectories to understand spreading risk in human crowds

### v1

@@ Abstract
A spreading process can be observed when particular information, substances, or diseases spread through a population over time in social and biological systems. It is widely believed that contact interactions among individual entities play an essential role in the spreading process. Although contact interactions are often influenced by geometrical conditions, little attention has been paid to understand their effects, especially on contact duration among pedestrians. To examine how the pedestrian flow setups affect contact duration distribution, we have analyzed trajectories of pedestrians in contact interactions collected from pedestrian flow experiments of uni-, bi- and multi-directional setups. Based on turning angle entropy and efficiency, we have classified the type of motion observed in the contact interactions. We have found that the majority of contact interactions in the unidirectional flow setup can be categorized as confined motion, hinting at the possibility of long-lived contact duration and hence high spreading risk. However, ballistic motion is more frequently observed in the other flow conditions, yielding frequent, brief contact interactions. Our results demonstrate that observing more confined motions is likely associated with the increase of parallel contact interactions regardless of pedestrian flow setups. This study highlights that the confined motions tend to yield longer contact duration, potentially leading to higher risk of virus exposure in human face-to-face interactions. These results have important implications for crowd management in the context of minimizing spreading risk.
@@ 1 Introduction
In this work, we present an improved version of the pedestrian contact interaction trajectory analysis studied in Ref. [28]. While Ref. [28] uses one test statistic for contact interaction trajectory characterization, this study applies a set of test statistics in an attempt to systematically characterize contact interaction trajectories for different pedestrian flow conditions.
@@ 2.1 Generation of synthetic trajectories
$\lambda=(c,1]$ with $c=0.1$
@@ 2.3 Classification criteria
First, we evaluate the scaling behavior of average end-to-end distance $\left\langle L\right\rangle$ as a function of average trajectory length $\left\langle l\right\rangle$. By doing that, we can quantify the confined and ballistic motions from OU process and CRW model simulations, respectively. Previous studies [45, 46] suggested that one can classify a random walk motion as a ballistic motion if the slope of $\left\langle L\right\rangle$ the curve in a log–log plot is approximately equal to 1, i.e., $\left\langle L\right\rangle\sim\left\langle l\right\rangle$. In addition, it is also suggested that $\left\langle L\right\rangle$ the curve reaches a plateau in the case of confined motion. That is, the trajectory of contact interaction is trapped in a small area, thus the end-to-end distance $L$ does not increase further, even though the trajectory length $l$ increases. Figure 3 shows a log–log plot of average end-to-end distance $L$ as a function of average trajectory length $l$ for the CRW model and OU process simulations. Analogous to previous studies [22, 25], we estimate the slope of $\left\langle L\right\rangle$ curve by fitting a function $a\log(t)+b$ to the estimated $\log(\left\langle L\right\rangle)$ curve. We choose the cutoff value of the exponent as $c=0.05$. When the slope of $\left\langle L\right\rangle$ curve in a log–log plot is between $1.951-c=$ 1.95 and $2.051+c=$ 2.05, i.e., $a\in[1.95,2.05]$, we consider the corresponding trajectory shows ballistic motion. If the value of $a$ is near 0, i.e., $a\in[-0.05,0.05]$, the corresponding trajectory is classified as confined motion. In line with the scaling behavior of $\left\langle L\right\rangle$ presented in Fig. 3, we categorize synthetic trajectories as the ballistic motion if the trajectories are generated by Eq. (2) with $\kappa\geq 300$, and trajectories generated by Eq. (1) as the confined motion.
@@
In addition, we define sub-ballistic motion for the remainder of the parameter space, which does not belong to either ballistic or confined motions. Figure 6 summarizes the presented classification criteria.
@@ 3.1 Datasets
From the experimental data, we extracted pairs of individuals in contact and their relative motion trajectories. We consider a pair of individuals to be in contact when the two individuals are within a contact radius. The contact radius $r_{c}$ would depend on the form of transmission in question. In this paper, we assume a 2 m radius based on previous studies [12, 32, 33, 34]. For the analysis, we consider contact interaction trajectories with at least 0.5 seconds records, i.e., 8 data points for the bi-directional, 2-way, and 3-way crossing flows, and 12 data points for the uni-directional and 4-way crossing flow based on literature [21, 25]. Table 4 shows the basic statistics of the selected scenarios, including the number of individuals $N$, experiment period, and the number of contacts. In this study, we consider three types of contact interactions: parallel, head-on, and crossing contact interactions. When two individuals in contact are moving in the same direction, one can define the parallel contact interaction. A head-on contact interaction is observed when two individuals are moving toward opposite directions. We can see a crossing contact interaction from a pair of individuals in contact when one individual is moving perpendicular to the other’s moving direction. The number of contacts is given as the number of interacting individual pairs. In Appendix B, we present the basic descriptive statistics of the other experimental scenarios.
@@ 3.2 Classification results
Based on the classification criteria presented in section 2.3, we have classified motion types of contact interaction trajectories. The summary of the motion type classification results for the selected scenarios can be found in Fig. 9. We can observe that ballistic motion is more frequently observed in the presented scenarios of the bi-directional (scenario bi-b10) and 4-way crossing flow (scenario crossing-90-a10). In contrast, the majority of contact interactions in the unidirectional flow setup (scenario uni-05) are categorized as confined motion, hinting at the possibility of long-lived contact duration. This implies that the risk level of virus exposure in human face-to-face intersections is high. In addition, for other experiment scenarios including crossing-90-d08 (2-way crossing setup) and crossing-120-b01 (3-way crossing setup), the sub-ballistic motion is the most frequently observed motion type. A summary of motion type classification results for other experiment scenarios are given in Appendix C.
@@ 4 Conclusion
A few selected experimental setups have been analyzed to study the fundamental role of pedestrian flow setups (e.g., uni-, bi-, and multi-directional flow) in the distribution of pedestrian motion types and contact duration. To generalize the findings of this study, the presented analysis should be further performed with larger number of scenarios and various layouts of pedestrian facilities at different locations. Another interesting extension of the presented study can be planned in line with the importance analysis of factors influencing on the motion types and duration of contact interactions [22, 27, 49, 50, 51].
@@ Acknowledgements
This research is supported by Ministry of Education (MOE) Singapore under its Academic Research Fund Tier 1 Program Grant No. RG12/21 MoE Tier 1.
@@ Appendix A Data and code availability
The data used for our study can be found from https://ped.fz-juelich.de/da/doku.php?id=start#data_section (last accessed 3 January, 2024). Table 5 shows short URL addresses where readers can find the descriptions and trajectory files of the presented experiment setups.
Table 5: Short URL addresses of the presented experiment setups (last accessed 3 January, 2024).

### v2

@@ Abstract
A spreading process can be observed when particular information, substances, or diseases spread through a population over time in social and biological systems. It is widely believed that contact interactions among individual entities play an essential role in the spreading process. Although contact interactions are often influenced by geometrical conditions, little attention has been paid to understand their effects, especially on contact duration among pedestrians. To examine how the pedestrian flow setups affect contact duration distribution, we have analyzed trajectories of pedestrians in contact interactions collected from pedestrian flow experiments of uni-, bi- and multi-directional setups. Based on turning angle entropy and efficiency, we have classified the type of motion observed in the contact interactions. We have found that the majority of contact interactions in the unidirectional flow setup can be categorized as confined motion, hinting at the possibility of long-lived contact duration. However, ballistic motion is more frequently observed in the other flow conditions, yielding frequent, brief contact interactions. Our results demonstrate that observing more confined motions is likely associated with the increase of parallel contact interactions regardless of pedestrian flow setups. This study highlights that the confined motions tend to yield longer contact duration, suggesting that the infectious disease transmission risk would be considerable even for low transmissibility. These results have important implications for crowd management in the context of minimizing spreading risk.
@@ 1 Introduction
In this work, we present an improved version of the pedestrian contact interaction trajectory analysis studied in Ref. [28]. To capture a wide range of contact interactions, we have generated synthetic trajectories based on random walk models whose characteristics are well understood. While Ref. [28] uses one test statistic for contact interaction trajectory characterization, this study applies a set of test statistics in an attempt to systematically characterize contact interaction trajectories for different pedestrian flow conditions. According to the previous studies [20, 21, 22], we have quantified the classification criteria for the contact interaction trajectories.
@@ 2.1 Generation of synthetic trajectories
$\lambda=(\lambda_{c},1]$ with $\lambda_{c}=0.1$
@@ 2.3 Classification criteria
First, we evaluate the scaling behavior of average end-to-end distance $\left\langle L\right\rangle$ as a function of average trajectory length $\left\langle l\right\rangle$. By doing that, we can quantify the confined and ballistic motions from OU process and CRW model simulations, respectively. Previous studies [48, 49] suggested that one can classify a random walk motion as a ballistic motion if the slope of $\left\langle L\right\rangle$ the curve in a log–log plot is approximately equal to 1, i.e., $\left\langle L\right\rangle\sim\left\langle l\right\rangle$. In addition, it is also suggested that $\left\langle L\right\rangle$ the curve reaches a plateau (i.e., the slope of $\log(\left\langle L\right\rangle)$ curve is near 0) in the case of confined motion. That is, the trajectory of contact interaction is trapped in a small area, thus the end-to-end distance $L$ does not increase further, even though the trajectory length $l$ increases. Figure 3 shows a log–log plot of average end-to-end distance $L$ as a function of average trajectory length $l$ for the CRW model and OU process simulations. Analogous to previous studies [22, 25], we estimate the slope of $\left\langle L\right\rangle$ curve by fitting a function $a\log(t)+b$ to the estimated $\log(\left\langle L\right\rangle)$ curve. We choose the cutoff value of the exponent as $c=0.05$. When the slope of $\left\langle L\right\rangle$ curve in a log–log plot is between $1.951-c=$ 1.95 and $2.051+c=$ 2.05, i.e., $a\in[1.95,2.05]$, we consider the corresponding trajectory shows ballistic motion. If the value of $a$ is between $-c=-0.05$ and $c=0.05$, i.e., $a\in[-0.05,0.05]$, the corresponding trajectory is classified as confined motion. In line with the scaling behavior of $\left\langle L\right\rangle$ presented in Fig. 3, we categorize synthetic trajectories as the ballistic motion if the trajectories are generated by Eq. (2) with $\kappa\geq 300$, and trajectories generated by Eq. (1) as the confined motion.
@@
In addition, we define sub-ballistic motion for the remainder of the parameter space, which does not belong to either ballistic or confined motions. Figure 6 summarizes the presented classification criteria. It is noted that change in the cutoff value $c$ affects the threshold values of $H$ and $E$ shown in Eqs. (13) and (14), but the scaling behavior of average end-to-end distance shown in Fig. 3 does not directly influence on the computation of $H$ and $E$ shown in Eqs. (7) and (8).
@@ 3.1 Datasets
From the experimental data, we extracted pairs of individuals in contact and their relative motion trajectories. We consider a pair of individuals to be in contact when the two individuals are within a contact radius. The contact radius $r_{c}$ would depend on the form of transmission in question, and is thus relevant to the evaluation of direct transmission risk. We have chosen a 2m radius based on existing literature [12, 32, 33], so that we can capture a wide range of common infectious diseases. It is also important to note that while we focus on spatial proximity as a measure of risk, there are many other factors that influence the risk of transmission (e.g., head orientations, respiratory activities, and airflow conditions) and that is highly disease dependent [34, 35, 36, 37, 38]. For the analysis, we consider contact interaction trajectories with at least 0.5 seconds records, i.e., 8 data points for the bi-directional, 2-way, and 3-way crossing flows, and 12 data points for the uni-directional and 4-way crossing flow based on literature [21, 25]. Table 4 shows the basic statistics of the selected scenarios, including the number of individuals $N$, experiment period, and the number of contacts. In this study, we consider three types of contact interactions: parallel, head-on, and crossing contact interactions. When two individuals in contact are moving in the same direction, one can define the parallel contact interaction. A head-on contact interaction is observed when two individuals are moving toward opposite directions. We can see a crossing contact interaction from a pair of individuals in contact when one individual is moving perpendicular to the other’s moving direction. The number of contacts is given as the number of interacting individual pairs. In Appendix B, we present the basic descriptive statistics of the other experimental scenarios.
@@ 3.2 Classification results
Based on the classification criteria presented in section 2.3, we have classified motion types of contact interaction trajectories. The summary of the motion type classification results for the selected scenarios can be found in Fig. 9. We can observe that ballistic motion is more frequently observed in the presented scenarios of the bi-directional (scenario bi-b10) and 4-way crossing flow (scenario crossing-90-a10). In contrast, the majority of contact interactions in the unidirectional flow setup (scenario uni-05) are categorized as confined motion, hinting at the possibility of long-lived contact duration. Such long-lived contact durations provide a longer time in which transmission can occur between a pair of individuals. However, in the case where individuals experience many long-lived interactions, it necessarily means fewer total interactions (i.e., a few long-term interactions) because an individual may have only a limited number of neighbors within a contact radius. The risk of the transmission then depends on the infectiousness of the disease, with long-lived interactions providing opportunity for transmission even in low transmissibility cases. This is in line with the findings of previous studies [12, 50], demonstrating that the infectious disease transmission risk would be considerable even for low transmissibility diseases. In addition, for other experiment scenarios including crossing-90-d08 (2-way crossing setup) and crossing-120-b01 (3-way crossing setup), the sub-ballistic motion is the most frequently observed motion type. A summary of motion type classification results for other experiment scenarios are given in Appendix C.
@@ 4 Conclusion
A few selected experimental setups have been analyzed to study the fundamental role of pedestrian flow setups (e.g., uni-, bi-, and multi-directional flow) in the distribution of pedestrian motion types and contact duration. To generalize the findings of this study, the presented analysis should be further performed with larger number of scenarios and various layouts of pedestrian facilities at different locations, for instance, mass gathering events like concerts and sport games. The presented approach can be applied to support facility managers and decision makers by quantifying the contact interaction characteristics for the crowd movement trajectories. Another interesting extension of the presented study can be planned in line with the importance analysis of factors influencing on the motion types and duration of contact interactions [22, 27, 51, 52, 53].
@@ Acknowledgements
This research is supported by Ministry of Education (MOE) Singapore under its Academic Research Fund Tier 1 Program Grant No. RG12/21 MoE Tier 1 and by the Singapore Ministry of Health’s National Medical Research Council under its National Epidemic Preparedness and Response R&D Funding Initiative (MOH-001041) Programme for Research in Epidemic Preparedness And REsponse (PREPARE).
@@ Appendix A Data and code availability
The data used for our study can be found from https://ped.fz-juelich.de/da/doku.php?id=start#data_section (last accessed 14 May, 2024). Table 5 shows short URL addresses where readers can find the descriptions and trajectory files of the presented experiment setups.
Table 5: Short URL addresses of the presented experiment setups (last accessed 14 May, 2024).